OpenStudy (anonymous):

log8+log3=logx-log10 findx

OpenStudy (anonymous):

Use the log formula: \[\large \log(a) + \log(b) = \log(a \times b)\] \[\large \log(a) - \log(b) = \log(\frac{a}{b})\]

OpenStudy (anonymous):

Can you solve it now??

OpenStudy (anonymous):

Are you getting or not??

OpenStudy (anonymous):

aswar?

OpenStudy (anonymous):

Sorry, I can't tell me I have given the formulas.. Just put them and show me your work..

OpenStudy (anonymous):

*you in place of me